Solve 2x2y – 11xy + 20y = 0.
Answer #1

Given differential equation 2x®y, – 11xy +20y = 0 Let y=x* d = mx*-7, y, = m(m-1)x*-? 2x*y- 11xy +20y = 0 2x* (m(m– 1)

answered by: ANURANJAN SARSAM